An ordinary perpetuity of $1000 per year offers a 7% return (i.e., discount rate is 7%). What is the future value of this perpetuity at the end of year 4? Round your answer to the nearest cent; do not include the $ sign (i.e., if the answer is $8,300.3562, enter it as 8,300.36).
Answer #1

Future Value = Periodic Payment*\frac{(1+r)^{n}-1}{r}

r = 7% OR 0.07

n = 4

= 1000 *\frac{(1+0.07)^{4}-1}{0.07}

= 1000 * 4.439943

= $ 4439.94